Elan  Pharmaceuticals, a  U.S.-based multinational pharmaceutical  company is evaluating an export sale of its  cholesterol-reduction drug with a prospective Indonesian distributor. The purchase would be for 1,650million Indonesian rupiah  (Rp). which at the current spot exchange rate of Rp9,470/$, translates into $174,234.42. Although not a big sale by company  standards company policy dictates that sales must be settled for at least a minimum gross  margin, in this  case, a cash settlement of $167,000. The current  90-day forward rate is Rp9,930/$. Although this rate appeared unattractive. Elan had to contact several major banks before even finding a forward quote on the rupiah. The consensus of currency forecasters at the  moment is that the rupiah will continue to  strengthen, possibly rising to Rp10,170/$ over the coming 90 to 120 days. Analyze the prospective sale and make a hedging recommendation.

1-How much in U.S. dollars will Elan receive in 90 days without a hedge if the expected spot rate in 90 days is assumed to be Rp9,470/$

2- How much in U.S. dollars will Elan receive in 90 days without a hedge if the expected spot rate in 90 days is assumed to be Rp10,170/$  

3- How much in U.S. dollars will Elan receive in 90 days without a hedge if the expected spot rate in 90 days is assumed to be Rp9930/$

4- How much in U.S. dollars will Elan receive in 90 days with a forward market  hedge?

5- Elan will have to decide whether making the sale into this specific market is worth breaking a company policy on minimum gross margin or taking significant currency risk by not using a forward  cover, The statement above is True or False ?
Answer #1

1). Amount received in 90-days if 90-day spot rate is Rp 9,470/$ = 1,650,000,000/9,470 = $174,234.42

2). Amount received in 90-days if 90-day spot rate is Rp 10,170/$ = 1,650,000,000/10,170 = $162,241.89

3). Amount received in 90-days if 90-day spot rate is Rp 9,930/$ = 1,650,000,000/9,930 = $166,163.14

4). Forward market hedge:

Sell forward at the current 90-day forward rate of Rp 9,930/$, then they will get

1,650,000,000/9,930 = $166,163.14 in 90 days.

5). True - If the company does not forward hedge then there is currency risk but at the given forward rate of Rp 9,930/$, it will be breaking its company policy of minimum gross margin.